# Cognitive-Mental-LLM: Evaluating Reasoning in Large Language Models for Mental Health Prediction via Online Text

### v2

@@ Abstract
Index Terms:
Large Language Models, Reasoning, Mental Health Prediction, Chain-of-Thought, Self-consistency, Tree-of-thought, Few-shot learning, Natural Language Processing, Online Text Analysis
@@ I Introduction
Traditional NLP-based classifiers, such as BERT [5] and RoBERTa [6], have demonstrated strong performance in general text classification but often struggle with mental health data due to subtle linguistic cues, contextual ambiguity, and the need for structured reasoning. Prior studies [7, 8, 9] have explored deep learning and transformer-based approaches, yet these models exhibit limitations in interpretability and robustness when applied to real-world mental health assessments.
@@ III Methodology
This study evaluates the effectiveness of structured reasoning techniques for mental health text classification using OpenAI’s o3-mini[21] model, a compact, reasoning-oriented language model known for its strong performance in scientific, mathematical, and programming tasks. We examine four prompting strategies—Chain-of-Thought (CoT), Tree-of-Thought (ToT), Few-shot CoT, and Self-Consistency (SC-CoT)—each enabling the model to generate interpretable reasoning prior to classification. This approach enhances both robustness and accuracy in mental health assessments. For comparison, we include results from fine-tuned models documented in a previous[7] study; however, no fine-tuning was conducted here, as our primary focus lies in evaluating the impact of different reasoning strategies.
@@ IV-A Classification Accuracy
Dataset CoT SC-CoT FS-CoT ToT Z-Shot* $\Delta$Z-Shot BERT*† $\Delta$BERT Alpaca*† FLAN-T5*† RoBERTa*†

### v3

@@ I Introduction
Traditional NLP-based classifiers, such as BERT [4] and RoBERTa [8], have demonstrated strong performance in general text classification but often struggle with mental health data due to subtle linguistic cues, contextual ambiguity, and the need for structured reasoning [7]. Prior studies [22, 18, 5] have explored deep learning and transformer-based approaches, yet these models exhibit limitations in interpretability and robustness when applied to real-world mental health assessments.
@@ III Methodology
x This study evaluates the effectiveness of structured reasoning techniques for mental health text classification using OpenAI’s o3-mini[11] model, a compact, reasoning-oriented language model known for its strong performance in scientific, mathematical, and programming tasks. We examine four prompting strategies—Chain-of-Thought (CoT), Tree-of-Thought (ToT), Few-shot CoT, and Self-Consistency (SC-CoT)—each enabling the model to generate interpretable reasoning prior to classification. This approach enhances both robustness and accuracy in mental health assessments. For comparison, we include results from fine-tuned models documented in a previous[22] study; however, no fine-tuning was conducted here, as our primary focus lies in evaluating the impact of different reasoning strategies.
@@ IV-A Classification Accuracy
[…]